OpenStudy (anonymous):

Can you divide 10/11 and 23/22. or is it undefined? I think it's undefined but I'm not sure.

OpenStudy (anonymous):

you can divide! Its result on periodically decimates.

OpenStudy (anonymous):

10/11/ 23/22 = 10/11 * 22/23 = 20/23

OpenStudy (anonymous):

when you talk about define and undefined you need to say what set of numbers you talking about.. real, natural, complex...

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

thats divisions is undefined on Natural numbers! =)